16.   Fifty-eight percent of all adults plan to purchase a Fitbit next year. You randomly select 200 adults and ask them if they plan to buy a Fitbit next year. What is the probability that at least 120 of these adults will say they will buy a Fitbit next year?
Answer #1

Mean = np = 0.58 * 200 = 116

Standard deviation = sqrt( np(1-p))

= sqrt ( 200 * 0.58 * 0.42)

= 6.98

Using normal approximation ,

P( X >= x) = P( Z > x-0.5 - mean / SD)

So,

P( X >= 120) = P( Z > 119.5 - 116 / 6.98)

= P( Z > 0.5014)

= 0.3080